Suppose the working-age population of a fictional economy falls into the following categories: 60 have full-time employment; 20 have part-time employment; 20 do not have employment, but are actively looking for employment; and 10 people are not employed and are not actively looking for employment. {Show all the steps including the relevant formulae & proper units}

  • (1) What is the size of Labor force
  • (2) What is the Unemployment rate
Answer #1

1. Ans: Size of Labor force = 100

Explanation:

Size of Labor force = Number of employed + Number of unemployed

= ( Full-time employment + Part-time employment ) + (Not employed , but are actively looking for employment)

= 60 + 20 + 20

= 100

2. Ans: Unemployment rate = 20%

Explanation:

Unemployment rate = (Total number of unemployed / Total labor force )* 100

= ( 20 / 100 ) *100

= 0.2 *100 = 20%